# Deep Multi-Threshold Spiking-UNet for Image Processing

### v2

@@ 4.4.2 Effect of Timesteps
Table 5 presents the performance with different timesteps on the CamSeq01 and BSD datasets. For the CamSeq01 dataset, mIoU gets better up to 20 steps and then it stays around 0.651. For the BSD dataset, the PSNR and SSIM get better as the steps go up. After 10 steps, these scores don’t change much and stay around 28.72 for PSNR and 0.802 for SSIM.
@@ 4.4.3 Energy Efficiency
Energy consumption is a critical factor when evaluating the cost of model inference. To provide a comprehensive assessment of the energy consumption of the Spiking-UNet model, we consider the energy cost of each operation and memory access based on the CamSeq01 dataset. The energy cost for Multiply-Accumulate (MAC) operations is estimated to be 4.6pJ, while Arithmetic and Comparison (AC) operations have an energy cost of 0.9pJ, as suggested by Horowitz et al. [49]. We utilize the method [50] to calculate the energy cost of accessing SRAM memory. Following the method described in [51], we define the FLOPs used to compute the energy consumption of the Spiking-UNet as
@@ 5 Conclusion and Discussion
Our research still has several limitations. As a preliminary exploration, we evaluate our Spiking-UNet on traditional and relatively small datasets for quick evaluation. In the future, Spiking-UNet will be tested on more newer and larger datasets. In addition, we only utilize Spiking-UNet on two image processing tasks. We will extend the application of Spiking-UNet, such as image super-resolution. What’s more, we will explore the deployment of Spiking-UNet on neuromorphic chips to validate its effectiveness in real-world.
@@ 6 Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grant 62032006 and 61901435.

### v4

@@ 4.4.2 Effect of Timesteps
Table 5 presents the performance with different timesteps on the CamSeq01 and BSD datasets. For the CamSeq01 dataset, mIoU gets better up to 20 steps and then it stays around 0.651. For the BSD dataset, the PSNR and SSIM get better as the steps go up. After 10 steps, these scores do not change much and stay around 28.72 for PSNR and 0.802 for SSIM.
@@ 4.4.3 Energy Efficiency
Energy consumption is a critical factor when evaluating the cost of model inference. To provide a comprehensive assessment of the energy consumption of the Spiking-UNet model, we consider the energy cost of each operation and memory access based on the CamSeq01 dataset. The energy cost for Multiply-Accumulate (MAC) operations is estimated to be 4.6pJ, while Arithmetic and Comparison (AC) operations have an energy cost of 0.9pJ, as suggested by Horowitz et al. horowitz20141 . We utilize the method lemaire2022analytical to calculate the energy cost of accessing SRAM memory. Following the method described in qu2023spiking , we define the FLOPs used to compute the energy consumption of the Spiking-UNet as
@@ 5 Conclusion and Discussion
Our research still has several limitations. As a preliminary exploration, we evaluate our Spiking-UNet on traditional and relatively small datasets for quick evaluation. In the future, Spiking-UNet will be tested on more newer and larger datasets. In addition, we only utilize Spiking-UNet on two image processing tasks. We will extend the application of Spiking-UNet, such as image super-resolution. Furthermore, we will explore the deployment of Spiking-UNet on neuromorphic chips to validate its effectiveness in real world.
@@ 6 Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grant 62021001.